OpenStudy (anonymous):

Exactly five athletes are competing to be in the starting lineup on a sports team, which consists of exactly five positions.If each player can play any position, and no one player will start at more than one position, how many different assignments of players to position are possible?

OpenStudy (anonymous):

\[5!=5\times 4\times 3\times 2\]